OpenStudy (anonymous):

3. What is the average (arithmetic mean) of all the multiples of ten from 10 to 190 inclusive? A. 90 B. 95 C. 100 D. 105 E. 110

OpenStudy (yolo_king):

C. 100

OpenStudy (anonymous):

Ok... Um thanks, can you help explain it to me?

OpenStudy (anonymous):

\(\frak\huge\color{#8a2be2}{Cant~give~direct~answers...}\)

OpenStudy (anonymous):

@iGreen

OpenStudy (yolo_king):

So what you do is that you got 10,20,30,40..190

OpenStudy (yolo_king):

To calcuate it you pretty much do 190=10+(n-1)10 n=19

OpenStudy (yolo_king):

s= (19/2)[a+1] =(19/2)*200/2 =1900

OpenStudy (yolo_king):

All you got to do is average it next and by doing that you need to divide it by 1900 by 19 to find 100

OpenStudy (anonymous):

O ok thank you very much but does arithmetic mean anything special? or is it another word for mean.

OpenStudy (anonymous):

yeah thats what it means

OpenStudy (anonymous):

O thx

OpenStudy (yolo_king):

Its simply the mean or the average of the sum of collection of numbers divided by the number of the collection of numbers.